A projectile is launched with an initial speed of 30 m/s and at an angle of 60 degrees with the horizontal.

a.) Find the horizontal and vertical components of its initial velocity
1
Expert's answer
2020-10-30T13:16:51-0400
"v_y=v\cos{60}=30\cos{60}=15\frac{m}{s}\\v_x=v\sin{60}=30\sin{60}=26\frac{m}{s}"